# Splitting Property of quasitriangular Hopf algebras

### v1

@@ 3. Twisted Tensor Product
If the Hopf algebras $K_{1},K_{2}$ above are not isomorphic to $𝕜\Bbbk$, we say that $(H,R)$ is a non-trivial twisted tensor product of them.
@@ 4. Main Results
    There exist normal left coideal subalgebras $L_{1},L_{2}$ of $(H,R)$ such that $H=L_{1}L_{2}$ is an exact factorization.
@@
(1$\Rightarrow$(2): We assume that $H=((K_{1}\otimes K_{2})^{J},R_{K_{1}\otimes K_{2}}^{J})$. Let $L_{1}=K_{1}\otimes 1\subseteq H$ and $L_{2}=S(1\otimes K_{2})\subseteq H$. It is clear that $K_{1}\otimes 1$ is normal left coideal subalgebra and $1\otimes K_{2}$ is normal right coideal subalgebra of $(H,R)$, i.e. a right coideal subalgebra such that $S(h_{1})kh_{2}\in K_{2}$ for all $h\in H,k\in K_{2}$. This implies that $L_{1},L_{2}$ are normal left coideal subalgebras of $(H,R)$. Now we see that the map $\varphi$ is bijective. Thus, the normal left coideal subalgebras $L_{1},L_{2}$ are obtained as required.
@@
By [20, Lemma 4.2], we see that $J$ is a twist for the quasitriangular Hopf algebra $K_{1}\otimes K_{2}$. We define $F:H\to K_{1}\otimes K_{2}$ as $F=(\pi_{1}\otimes\pi_{2})\circ\Delta$. Then $F$ is a Hopf map. Moreover, $(F\otimes F)(R)=R_{K_{1}\otimes K_{2}}^{J}$, where
@@
By Theorem 4.2, we only need to show that there exist normal left coideal subalgebras $L_{1},L_{2}$ such that the following map is bijective
@@
Let $L_{1}=H^{co\pi}$ and let $L_{2}=\Phi_{R}(K^{*})$ (See Equation (3.1) for the definition). We obtain that $L_{1},L_{2}$ are normal left coideal subalgebras of $(H,R)$. To show $\varphi$ is bijective, we only need to show that $\varphi=\varphi^{\prime}$ as linear maps via identifying $\underline{H}$ and $H$ as vector spaces. By Lemma 2.4, we get $\underline{H^{co\pi}}=H^{co\pi}$ as vector spaces. Since the definition of $\gamma$, we know $\operatorname{im}(\gamma)=\Phi_{R}(K^{*})$ as vector spaces. Thus, the map $\varphi$ is just $\varphi^{\prime}$ as linear maps, which implies that $\varphi$ is bijective. Hence by Theorem 4.2, $(H,R)$ is a twisted tensor product of $(K,\pi)$ and $(K^{\prime},\pi^{\prime})$ for some quotient Hopf algebra $(K^{\prime},\pi^{\prime})$ of $(H,R)$. ∎
@@
Note that $L_{1}$ is left normal coideal subalgebra. We only have to show $L_{2}$ is also a left normal coideal subalgebra. Recall that $L_{2}=l_{R}(K^{*})$ is Hopf subalgebra of $H$, we have to prove that $L_{2}$ is normal. Since $\varphi$ is bijective, we get $H=L_{1}L_{2}$. Next, we show $(ab)_{(1)}cS((ab)_{(2)})\in L_{2}$ for all $a\in L_{1}$ and $b,c\in L_{2}$. Since $L_{2}$ is Hopf subalgebra, we obtain that $b_{(1)}cS(b_{(2)})\in L_{2}$. Thus, we can assume $b_{(1)}cS(b_{(2)})=l_{R}(f\circ\pi)$ for some $f\in K^{*}$. Since
@@
$a_{(1)}l_{R}(f\circ\pi)S(a_{(2)})=l_{R}(f\circ\pi)\epsilon(a).$
Thus, we have shown that $L_{2}$ is normal.

### v2

@@ 3. Twisted Tensor Product
If the Hopf algebras $K_{1},K_{2}$ above are not isomorphic to $𝕜\Bbbk$, we say that $(H,R)$ is a non-trivial twisted tensor product of them. Note that the Hopf algebra $(H,R)$ in this definition may be triangular, and we will consider some of these cases (eg. Theorem 4.10).
@@ 4. Main Results
    There exist normal left coideal subalgebras $L_{1},L_{2}$ of $(H,R)$ such that
    • $H=L_{1}L_{2}$ is an exact factorization;
    • $(\pi_{1}\otimes\pi_{2})(R_{21}R)=1\otimes 1$, where $\pi_{1}$(resp. $\pi_{2}$) denote the associated natural quotient map with $H/HL_{1}^{+}$(resp. $H/HL_{2}^{+}$);
@@
(1$\Rightarrow$(2): We assume that $H=((K_{1}\otimes K_{2})^{J},R_{K_{1}\otimes K_{2}}^{J})$. Let $L_{1}=K_{1}\otimes 1\subseteq H$ and $L_{2}=S(1\otimes K_{2})\subseteq H$. It is clear that $K_{1}\otimes 1$ is normal left coideal subalgebra and $1\otimes K_{2}$ is normal right coideal subalgebra of $(H,R)$, i.e. a right coideal subalgebra such that $S(h_{1})kh_{2}\in K_{2}$ for all $h\in H,k\in K_{2}$. This implies that $L_{1},L_{2}$ are normal left coideal subalgebras of $(H,R)$. Moreover, $(\pi_{1}\otimes\pi_{2})(R_{21}R)=1\otimes 1$ in this case. Now we see that the map $\varphi$ is bijective. Thus, the normal left coideal subalgebras $L_{1},L_{2}$ are obtained as required.
@@
By [20, Lemma 4.2], we see that $J$ is a twist for the quasitriangular Hopf algebra $K_{1}\otimes K_{2}$. We define $F:H\to K_{1}\otimes K_{2}$ as $F=(\pi_{1}\otimes\pi_{2})\circ\Delta$. Then $F$ is a Hopf map. Moreover, $(F\otimes F)(R)=R_{K_{1}\otimes K_{2}}^{J}$ by using $(\pi_{1}\otimes\pi_{2})(R_{21}R)=1\otimes 1$, where
@@
By Theorem 4.2, we only need to show that there exist normal left coideal subalgebras $L_{1},L_{2}$ such that the following map is bijective and they satisfy $(\pi_{1}\otimes\pi_{2})(R_{21}R)=(1\otimes 1)$, where $\pi_{1}$(resp. $\pi_{2}$) denote the associated natural quotient map with $H/HL_{1}^{+}$(resp. $H/HL_{2}^{+}$)
@@
Let $L_{1}=H^{co\pi}$ and let $L_{2}=\Phi_{R}(K^{*})$ (See Equation (3.1) for the definition). We obtain that $L_{1},L_{2}$ are normal left coideal subalgebras of $(H,R)$. To show $\varphi$ is bijective, we only need to show that $\varphi=\varphi^{\prime}$ as linear maps via identifying $\underline{H}$ and $H$ as vector spaces. By Lemma 2.4, we get $\underline{H^{co\pi}}=H^{co\pi}$ as vector spaces. Since the definition of $\gamma$, we know $\operatorname{im}(\gamma)=\Phi_{R}(K^{*})$ as vector spaces. Thus, the map $\varphi$ is just $\varphi^{\prime}$ as linear maps, which implies that $\varphi$ is bijective. Since $L_{2}=\Phi_{R}(K^{*})$ and the definition of $\pi_{1}$, we have $\pi_{1}(\Phi_{R}(f))=\epsilon(f)1$ for any $f\in K^{*}$, this implies $(\pi_{1}\otimes\pi_{2})(R_{21}R)=(1\otimes 1)$. Hence by Theorem 4.2, $(H,R)$ is a twisted tensor product of $(K,\pi)$ and $(K^{\prime},\pi^{\prime})$ for some quotient Hopf algebra $(K^{\prime},\pi^{\prime})$ of $(H,R)$. ∎
@@
Note that $L_{1}$ is left normal coideal subalgebra. We only have to show $L_{2}$ is also a left normal coideal subalgebra and $(\pi_{1}\otimes\pi_{2})(R_{21}R)=(1\otimes 1)$, where $\pi_{1}$(resp. $\pi_{2}$) denote the associated natural quotient map with $H/HL_{1}^{+}$(resp. $H/HL_{2}^{+}$). Recall that $L_{2}=l_{R}(K^{*})$ is Hopf subalgebra of $H$, we have to prove that $L_{2}$ is normal. Since $\varphi$ is bijective, we get $H=L_{1}L_{2}$. Next, we show $(ab)_{(1)}cS((ab)_{(2)})\in L_{2}$ for all $a\in L_{1}$ and $b,c\in L_{2}$. Since $L_{2}$ is Hopf subalgebra, we obtain that $b_{(1)}cS(b_{(2)})\in L_{2}$. Thus, we can assume $b_{(1)}cS(b_{(2)})=l_{R}(f\circ\pi)$ for some $f\in K^{*}$. Since
@@
(4.4) $a_{(1)}l_{R}(f\circ\pi)S(a_{(2)})=l_{R}(f\circ\pi)\epsilon(a).$
Thus, we have shown that $L_{2}$ is normal. Further, one can see that $H^{co\pi}$ commutes with $l_{R}(K^{*})$ due to above equality (4.4). Thus, $H=H^{co\pi}.l_{R}(K^{*})$ has actually tensor product structure as Hopf algebras. From this, we can assume that $H=A\otimes B$ as Hopf algebras and
$\pi_{1}=(\operatorname{Id}_{A}\otimes\epsilon_{B}),\;\pi_{2}=(\epsilon_{A}\otimes\operatorname{Id}_{B}),$
where $A=H^{co\pi}$ and $B=l_{R}(K^{*})$. From this, we can assume $R=\sum_{i,j}(a_{i}\otimes b_{j})\otimes(a^{i}\otimes b^{j})$, where $a_{i},a^{i}\in A$ and $b_{i},b^{i}\in B$. Then, we have:
$(\pi_{1}\otimes\pi_{2})(R_{21}R)$ $=(\pi_{1}\otimes\pi_{2})[\sum_{k,l}(a^{k}\otimes b^{l})\otimes(a_{k}\otimes b_{l})][\sum_{i,j}(a_{i}\otimes b_{j})\otimes(a^{i}\otimes b^{j})]$
$=\epsilon_{A}(a_{k})\epsilon_{B}(b_{j})\epsilon_{A}(a^{i})\epsilon_{B}(b^{l})(a^{k}a_{i}\otimes b_{l}b^{j})$
$=\epsilon_{A}(a^{i})\epsilon_{B}(b^{l})(a_{i}\otimes b_{l})$
$=1\otimes 1,$
where the last equality follows from $\sum_{i,l}(a_{i}\otimes b_{l})\otimes(\epsilon_{A}(a^{i})\otimes\epsilon_{B}(b^{l}))=1\otimes 1$ and the penultimate equality follows from an analogous equation. Thus, we have $(\pi_{1}\otimes\pi_{2})(R_{21}R)=(1\otimes 1)$.